OpenStudy (anonymous):

solve the difference equation √Sn-1+√Sn-2+√Sn-3+√.......... given that S0=4;

OpenStudy (anonymous):

isn't that supposed to be something like:\[S_n=\sqrt{S_{n-1}}+\sqrt{S_{n-2}}+\sqrt{S_{n-3}}+...\]

OpenStudy (anonymous):

if it is the rec formula then use this hint : \[S_n=\sqrt{S_{n-1}}+\sqrt{S_{n-2}}+\sqrt{S_{n-3}}+...=\sqrt{S_{n-1}}+S_{n-1}\]for n>1